Jason liked to rent cars from Airport Rent-a-car. The first time he rented a car, he drove 125 miles over 2 days. On his second trip, Jaosn drove 350 miles over 4 days. The cost of the first thrip was $95.75. The cost of the second trop was $226.50. What is the cost per day and cost per mile of renting a car from Airport Rent-a-car?
Answer by RAY100(1637)

Let D=day rate
Let M= Mileage rate
1) 2D +125M=95.75
2) 4D+350M=226.50
mult (1) by 2
4D +250M= 191.50
subtract from (2)
100M= 35.00
divide by 100 both sides of eqn
M= .35
subst in (1)
2D +(.35)125 = 95.75
2D + 43.75 = 95.75
subtract 43.75 both sides
2D= 52
divide by 2 both sides
D= 26

check
2(26) + .35 ( 125) = 95.75 ok
4(26) + .35 ( 350) = 226.5 ok